# Rotating Einstein-Maxwell-Dilaton Black Hole as a Particle Accelerator

### v1

@@ Abstract
Similar to particle accelerators, black holes also have the ability to accelerate particles, generating significant amounts of energy through particle collisions. In this study, we examine the horizon and spacetime structures of a rotating black hole within the framework of Einstein-Maxwell-Dilaton gravity. Additionally, we extend the analysis to explore particle collisions and energy extraction near this black hole using the Banados-Silk-West mechanism. Our findings reveal that the mass and angular momentum of the colliding particles significantly influence the center of mass energy, more so than the parameters of the black hole itself. Furthermore, we apply the Banados-Silk-West mechanism to massless particles, particularly photons, while disregarding their intrinsic spin in plasma; an aspect that has not been previously explored. The Banados-Silk-West mechanism cannot be directly applied, as the refractive index condition only permits photon propagation, meaning that massive particles in vacuum cannot be included in this study. We derive the propagation conditions for photons and analyze photon collisions by treating them as massive particles in a dispersive medium. The impact of the plasma parameter on the extracted center of mass energy is also examined. Our results show that the plasma parameter has a relatively weak and unchanged effect on the center of mass energy across all cases, indicating that energy losses due to friction within the medium are a contributing factor.
Keywords: Einstein-Maxwell-Dilaton black hole; Center of mass energy; Plasma
March 18, 2025
@@ I Introduction
A black hole can be conceptualized as a particle accelerator due to the immense gravitational forces near its event horizon. A significant amount of energy can be extracted from particle collisions occurring in the strong gravitational field around black holes. One might question the existence of particles within the black hole’s field; however, due to the influence of gravity and accretion disks, matter likely exists in the vicinity of black holes, where numerous particle collisions can occur. The energy extracted from these particle collisions is typically measured in terms of center of mass energy (CME) for the system, rather than the energies of the individual constituents. This is a purely relativistic phenomenon and plays a crucial role in high-energy astrophysics. Baushev [1] was the first to propose that a black hole could act as a particle accelerator. He examined the Schwarzschild black hole and discovered that particles near the black hole could collide, resulting in the production of a substantial amount of energy. However, his study did not provide a formal or detailed analysis of particle collisions in the vicinity of a black hole. A few months later, Bañados et al. [2] developed a formal procedure to calculate the energy generated by particle collisions near a black hole. They considered two uncharged particles of equal mass colliding near a Kerr black hole and found a significant amount of CME. They also extended the analysis to the case of a Schwarzschild black hole. This method later became known as the Banados-Silk-West (BSW) mechanism. However, Berti et al. [3] proposed that such collisions are less likely to occur in nature, by discussing the BSW mechanism. They found that the upper bound on the spin of the Kerr black hole, along with radiation losses, imposes an upper limit on the CME. Additionally, Jacobson and Sotiriou [4] highlighted the physical limitations of such high-energy collisions. In the same year, Lake [5, 6] suggested that classical black holes are internally unstable due to a divergent CME for two neutral, massive colliding particles at the Cauchy horizon of a Kerr black hole. Grib and Pavlov [7, 8, 9] investigated particle collisions and observed a high CME resulting from scattering in both non-extremal and extremal Kerr black holes. Harada and Kimura [10] derived an analytic, general relation to calculate the center-of-mass energy (CME) for the collision of two particles with equal masses near the event horizon of a Kerr black hole. This relation was applied to two distinct collision scenarios: in the first, a particle traveling from the innermost stable circular orbit collided with another particle closer to the event horizon, while in the second, both particles were in the innermost stable circular orbit. They identified the maximum CME limits for extremal Kerr black holes. Based on the BSW mechanism some notable works are also given in the Refs. [11, 12, 13, 14, 15, 16, 17, 18, 19] and some of the recent advances in particle collision theory and energy extraction from the center of mass frame by using the BSW mechanism is also given in the Refs. [20, 21, 22, 23].

### v2

@@ Abstract
Similar to particle accelerators, black holes also have the ability to accelerate particles, generating significant amounts of energy through particle collisions. In this study, we examine the horizon and spacetime structures of a rotating black hole within the framework of Einstein-Maxwell-Dilaton gravity. Additionally, we extend the analysis to explore particle collisions and energy extraction near this black hole using the Bañados-Silk-West mechanism. Our findings reveal that the mass and angular momentum of the colliding particles significantly influence the center of mass energy, more so than the parameters of the black hole itself. Furthermore, we apply the Bañados-Silk-West mechanism to massless particles, particularly photons, while disregarding their intrinsic spin in plasma; an aspect that has not been previously explored. The Bañados-Silk-West mechanism cannot be directly applied, as the refractive index condition only permits photon propagation, meaning that massive particles in vacuum cannot be included in this study. We derive the propagation conditions for photons and analyze photon collisions by treating them as massive particles in a dispersive medium. The impact of the plasma parameter on the extracted center of mass energy is also examined. Our results show that the plasma parameter has a relatively weak and unchanged effect on the center of mass energy across all cases, indicating that energy losses due to friction within the medium are a contributing factor.
July 22, 2025
@@ I Introduction
A black hole can be conceptualized as a particle accelerator due to the immense gravitational forces near its event horizon. A significant amount of energy can be extracted from particle collisions occurring in the strong gravitational field around black holes. One might question the existence of particles within the black hole’s field; however, due to the influence of gravity and accretion disks, matter likely exists in the vicinity of black holes, where numerous particle collisions can occur. The energy extracted from these particle collisions is typically measured in terms of the center of mass energy (CME) for the system, rather than the energies of the individual constituents. This is a purely relativistic phenomenon and plays a crucial role in high-energy astrophysics. When a black hole has rotation, the spacetime structure notably changes from the static case, creating a region known as the ergosphere, within which the phenomenon known as frame dragging becomes significant. This effect inevitably drags any particle or radiation within this region in the direction of the black hole’s rotation, introducing important additional dynamics for the analysis of particle collisions near these compact objects [1]. Moreover, considering modified gravity theories such as Einstein-Maxwell-Dilaton (EMD) gravity is particularly relevant since these theories naturally emerge from fundamental contexts like string theory or supergravity models. Specifically, the additional scalar field known as the dilaton not only modifies the structure of the black hole itself but also significantly influences how matter and radiation interact in its vicinity. This influence is critical because coupled scalar fields can substantially alter horizon formation, spacetime geometry near the black hole, and extreme energetic processes such as high-energy particle collisions [2, 3, 4]. Additionally, explicitly considering a dispersive medium, such as plasma, becomes essential since plasma is commonly found around real black holes. Thus, studying photon propagation and collisions in these dispersive media—effectively treated as massive particles due to plasma dispersive properties—allows modeling realistic conditions under which emitted radiation and observable processes in extreme astrophysical environments could be notably influenced. By investigating these phenomena through some mechanism (see below), this study deepens the theoretical understanding of particle dynamics in intense gravitational fields and offers quantitative predictions potentially testable by future high-precision astronomical observations, thereby strengthening the connection between theory and observation in relativistic astrophysics.
@@
Baushev [5] was the first to propose that a black hole could act as a particle accelerator. He examined the Schwarzschild black hole and discovered that particles near the black hole could collide, resulting in the production of a substantial amount of energy. However, his study did not provide a formal or detailed analysis of particle collisions in the vicinity of a black hole. Later on, Bañados et al. [6] developed a formal procedure to calculate the energy generated by particle collisions near a black hole. They considered two uncharged particles of equal mass colliding near a Kerr black hole and found a significant amount of CME. They also extended the analysis to the case of a Schwarzschild black hole. This method later became known as the Bañados-Silk-West (BSW) mechanism. However, Berti et al. [7] proposed that such collisions are less likely to occur in nature, by discussing the BSW mechanism. They found that the upper bound on the spin of the Kerr black hole, along with radiation losses, imposes an upper limit on the CME. Additionally, Jacobson and Sotiriou [8] highlighted the physical limitations of such high-energy collisions. In the same year, Lake [9, 10] suggested that classical black holes are internally unstable due to a divergent CME for two neutral, massive colliding particles at the Cauchy horizon of a Kerr black hole. Grib and Pavlov [11, 12, 13] investigated particle collisions and observed a high CME resulting from scattering in both non-extremal and extremal Kerr black holes. Harada and Kimura [14] derived an analytic, general relation to calculate the center-of-mass energy (CME) for the collision of two particles with equal masses near the event horizon of a Kerr black hole. This relation was applied to two distinct collision scenarios: in the first, a particle traveling from the innermost stable circular orbit collided with another particle closer to the event horizon, while in the second, both particles were in the innermost stable circular orbit. They identified the maximum CME limits for extremal Kerr black holes. Based on the BSW mechanism some notable works are also given in the Refs. [15, 16, 17, 18, 19, 20, 21, 22, 23] and some of the recent advances in particle collision theory and energy extraction from the center of mass frame by using the BSW mechanism is also given in the Refs. [24, 25, 26, 27].
@@ Acknowledgments
Conflict of Interest
The authors declare no conflict of interest.
Data Availability Statement
Data sharing is not applicable to this article as no new data were created or analyzed in this study.
Keywords
Einstein-Maxwell-Dilaton black hole; Center of mass energy; Plasma; Particle accelerator.